OpenStudy (anonymous):

I need some assistance with this question below. I have went through the practice one that walks me through it but still am missing something.

OpenStudy (anonymous):

\[A=p(1+\frac{ r }{ n })^{nt}\] p= 200, r= 0.07, n= 1, t= 3 (round to nearest hundredth) --Compound Interest Formula

OpenStudy (valpey):

\[A=200(1+\frac{0.07}{1})^{1*3}=200(1.07)^3\]

OpenStudy (anonymous):

Yes and they you distribute the 200 into the parentheses. which you have done but still was coming up with the wrong answer.

OpenStudy (valpey):

No. Keep the 200 outside of the parenthesis until after you have cubed 1.07

OpenStudy (anonymous):

So you would not distribute to get rid of the parenthesis to get rid of them? That does not make sense.

OpenStudy (valpey):

Exponents come before multiplication in order of operations.\[200(1.07)^3=200(1.07^3)=200*1.225\]

OpenStudy (anonymous):

So wouldn't you have to take your 1 to the exponent before you got to the 0.07/1?

OpenStudy (valpey):

No. Order of operations is Parentheses, Exponents, Multiplication/Division, Addition/Subtraction. First you simplify the inside of the parentheses, Second you reconcile exponents, third you multiply by 200.

OpenStudy (anonymous):

Okay so then it would have to come out as A=200(1.07)^3 then I would do 1.07^3=1.225 then multiply 200(1.225) = 245. Correct?

OpenStudy (anonymous):

Which that is what I had tried but it did not come out as correct.

OpenStudy (valpey):

Looks right.

OpenStudy (anonymous):

It did not come out though... I put in the answer I came out with but it said it was incorrect.

OpenStudy (valpey):

If you put $200 in an account earning 7% effective interest per year, after three years you would have $245.

OpenStudy (anonymous):

Right but it seem as though it is missing something because it looks like it has to be coming out as a decimal according to the instructions to round to the nearest hundredth place.

OpenStudy (valpey):

t=0 $200; t=1 $214; t=2 $228.98; t=3 $245.01;

OpenStudy (anonymous):

Now how did you come up with that?

OpenStudy (valpey):

Multiply 200 by 1.07 then do it again and again.

OpenStudy (valpey):

At the end of each year you have 1.07 times as much as you did at the beginning of the year (or 7% more).

OpenStudy (anonymous):

Even after trying to do 200(1.07)^3, when I round the answer to the nearest hundredth it still comes out as 245.10 but that still is not the answer... maybe this stupid "MyMathLap" online thing is wrong.

OpenStudy (shane_b):

You didn't correctly round to the nearest hundredth: \[A=p(1+\frac{ r }{ n })^{nt}=200(1+\frac{0.07}{1})^{1*3}=200(1.07)^3=$245.01\]

OpenStudy (anonymous):

Just miss typed the .01 as .10... got it thanks!

OpenStudy (shane_b):

np :)